An unknown gas occupies 10dm³ of space at 4atm and 293kelvin.what is the amount of moles of gas present if the gas iss a real gas? Take R=0.082atm/(dm³k)
1
Expert's answer
2020-03-09T09:44:05-0400

Clapeyron-Mendeleev equation:

"PV=nRT",

where P - pressure, V - volume, n - amount of the gas, R - gas constant, T - temperature.

"n=\frac{PV}{RT}"

"n=\frac{4 \cdot 10}{0.082 \cdot 293} \approx 1.66 \; mol"


Answer: 1.66 mol.